OpenStudy (anonymous):

k=1/3(b-2a) solve for b would you use the distributive property first or divide both sides by 1/3?

OpenStudy (nurali):

K =1/3 (b – 2a) 3K = (b – 2a) 3K + 2a = b b = 3K + 2a

OpenStudy (nurali):

if correct please click best response

OpenStudy (anonymous):

Why would you divide 1/3 before you would use the distributive property?

OpenStudy (anonymous):

since 1/3 is only on one side u can move it or divide and the result will bee the same 3k=b-2a and b=3k+2a

OpenStudy (anonymous):

\[k=\frac{ 1 }{ 3 }*(b-2a)| *3 <=> 3k=b-2a <=> b=3k+2a\]